# Attosecond physics hidden in Cherenkov radiation

### v2

@@ Abstract
Cherenkov radiation of charged particles moving with superluminal velocities in transparent media is a well-studied phenomenon with a plethora of applications. Its microscopic origins can be traced to the polarization of atomic shells, characterized by time scales in the subfemtosecond range — dynamics that eludes conventional macroscopic treatment. Here we present a theoretical framework for probing the intrinsic dynamics of Cherenkov radiation, unveiling quantum features absent in classical realm and even in a fully quantum theory in momentum space. These features include a finite formation length and spreading time of the photon, the latter becoming negative nearby the Cherenkov angle, a finite flash duration tied to the size of the electron packet, along with a shift in the photon arrival time that can be either positive or negative. The calculated time scales naturally lie in the attosecond range for relevant parameters, thereby linking this macroscopic phenomenon back to its atomic origins. Finally, we propose that by measuring the duration of the Cherenkov flash one can in principle retrieve the length of the emitting packet, deepening our understanding of quantum coherence effects in photon emission.
@@ Introduction
Cherenkov radiation (ChR) by charged particles in media [1, 2, 3, 4, 5, 6, 7, 8, 9, 10] is the simplest example of a wide range of phenomena embracing transition, diffraction, Smith-Purcell radiation, and other mechanisms of photon emission [11, 12, 13, 14]. Their common microscopic origin is atomic bremsstrahlung [15] due to dynamic polarization of atomic shells by the field of the charge, and the characteristic time scales are femto- and attoseconds, typical for the AC Stark effect [16]. Along with numerous applications of classical ChR in particle detection [17], neutrino telescopes [18, 19], in gamma-ray astronomy [20, 21], and other fields, it has also recently attracted attention of biomedical community as a new tool for molecular imaging and therapy in cancer treatment [22, 23, 24]. However, despite the growing interest to inherently quantum features of ChR [25, 26, 27], a solid link between microscopic and macroscopic theories remains elusive, and the atomic dynamics stays hidden even in the fully quantum treatment of ChR in momentum space.
@@
Here we point out that one can access atomic time scales in ChR and in its generalizations by using a quantum theory in phase space where – similar to quantum optics [28] – we employ a Wigner function to characterize the emitted photon field. We demonstrate how to probe the field in the pre-wave (or formation) zone [29, 12, 30] in which the partial waves interfere, spread, and the Cherenkov cone is not formed yet. Most strikingly, our theory establishes a link between the evolving coherence length of the electron packet, the Cherenkov flash duration, and a quantum temporal delay that the photon experiences in medium. This delay can be either positive or negative and belongs to the attosecond range, typical for atomic excitation processes studied by the attosecond spectroscopy and metrology [31]. The quantum shift in the photon arrival time, coupled with the finite flash duration and the spreading dynamics, unveils intricate coherence properties that enable refined temporal control in quantum emission processes, enriching the landscape of quantum optics, ultrafast physics, and of applications of ChR and related phenomena. The system of units $\hbar=c=1$ is used, the electron velocity is $\beta=u_{p}/c\equiv u_{p}$, $m=0.511$ MeV is the electron mass, $1/m\equiv\hbar/mc=3.86\times 10^{-11}$ cm, and $t_{c}=1/m\equiv\hbar/mc^{2}\approx 1.3\times 10^{-21}\,\text{s}$.
Figure 1: Generation of Cherenkov radiation by an electron packet with an rms size $\sigma_{x}$ in a dielectric medium with a refractive index $n$. The photon field gradually becomes separated from the electron in the formation zone and spreads. The Cherenkov cone is formed in the far-field only if the electron is not detected. The phase space picture implies that the photon with a momentum ${\bf k}$ is detected in a region of space centered at the point ${\bf r}$ at the moment of time $t$.
@@
When an electron emits a photon, the two-particle state within the first order of the perturbation theory is $|e^{\prime},\gamma\rangle=(\hat{1}+\hat{S}^{(1)})|\text{in}\rangle$, where $|\text{in}\rangle=|e_{\text{in}}\rangle\otimes|0_{\gamma}\rangle$ and $\hat{S}^{(1)}=-ie\int d^{4}x\,\hat{j}^{\mu}(x)\hat{A}_{\mu}(x)$ [32]. Fig.1 illustrates how the photon packet gradually becomes spatially separated from the electron, spreads, and eventually turns into a plane wave propagating at the angles $\theta,\phi$. If we detect the electron in a plane-wave state $\langle{\bf p}^{\prime},\lambda^{\prime}|$, $\lambda^{\prime}=\pm 1/2$, the state of the photon becomes
$\displaystyle|\gamma\rangle=\langle{\bf p}^{\prime},\lambda^{\prime}|e_{\text{in}}\rangle|0_{\gamma}\rangle+\sum\limits_{\lambda_{\gamma}}\int\frac{d^{3}k}{(2\pi)^{3}}\,|{\bf k},\lambda_{\gamma}\rangle\,S_{fi}.$ (1)
The second term here is a coherent superposition of plane waves with the momenta ${\bf k}$ and the helicity $\lambda_{\gamma}=\pm 1$ and a matrix element is $S_{fi}=\langle{\bf k},\lambda_{\gamma};{\bf p}^{\prime},\lambda^{\prime}|\hat{S}^{(1)}|\text{in}\rangle$. We treat the incoming electron as a Gaussian packet with a wave function $f_{e}^{(\text{in})}({\bf p},\lambda)=\langle{\bf p},\lambda|e_{\text{in}}\rangle$, a mean momentum $\langle{\bf p}\rangle$, an uncertainty $\sigma\ll m$, which means that the rms size – also called coherence length – of the packet is much larger than the Compton wavelength, $\sigma_{x}=1/\sigma\gg 1/m$. For the moment, we take a simplified model with the packet, spherically symmetric in the laboratory frame, $\sigma_{\perp}=\sigma_{||}\equiv\sigma_{x}=1/\sigma$.
Let us define the energy density of the photon field in real space and time, an electric part of which is (Sec.1,2 in Supplementary)
$\displaystyle\frac{{\displaystyle 1}}{{\displaystyle 8\pi}}\langle\gamma|\hat{\bm{E}}^{2}({\bm{r}},t)|\gamma\rangle\to\frac{{\displaystyle 1}}{{\displaystyle 4\pi}}\left|\langle 0|\hat{\bm{E}}({\bm{r}},t)|\gamma\rangle\right|^{2}=\int\frac{d^{3}k}{(2\pi)^{3}}\,\mathcal{W}({\bf r},{\bf k},t),$ (2)
where we have subtracted a contribution of the vacuum energy and $\hat{\bm{E}}({\bm{r}},t)$ is a secondary-quantized electric field operator [33]. A contribution of the magnetic field can be written similarly. Here
@@
is a Wigner function of the photon, $\bm{E}_{\lambda_{\gamma}}({\bf k})=\frac{i\omega\sqrt{4\pi}}{\sqrt{2\omega n^{2}}}\,{\bm{e}}_{{\bf k}\lambda_{\gamma}}\sum\limits_{\lambda}\int\frac{d^{3}p}{(2\pi)^{3}}\,f_{e}^{(\text{in})}({\bf p},\lambda)S_{fi}^{(\text{pw})},\,{\bm{e}}_{{\bf k}\lambda_{\gamma}}\cdot{\bf k}=0,$ is a positive-frequency component of its electric field.
According to the standard interpretation [33], the energy density (2) defines probability to detect a photon in a region of space centered at the point ${\bf r}$ at the moment of time $t$. Clearly, the second marginal distribution $\int d^{3}x\,\mathcal{W}({\bf r},{\bf k},t)\propto|S_{fi}^{(\text{pw})}|^{2}$ yields probability to detect a photon with the frequency $\omega$ and a wave vector ${\bf k},\,|{\bf k}|=n(\omega)\,\omega$, the standard result of the quantum theory in momentum space (see [26]). Therefore, it is this Wigner function (3) in phase space that contains all the information on spatial distribution of the photon field at a given distance ${\bf r}$ – also in the near-field zone – and on its dynamics. Consequently, emission takes place in the pre-wave zone even if the condition of ChR is not met, but the waves do not constructively interfere to form a cone in the far field.
We calculate the Wigner function in the paraxial approximation, $\sigma\ll m$, in a medium with weak dispersion, $\frac{\omega}{n(\omega)}\frac{dn(\omega)}{d\omega}\ll 1$, and represent the amplitude [32, 26] as follows:
$S_{fi}^{(\text{pw})}=|S_{fi}^{(\text{pw})}|\,e^{i\zeta_{fi}}.$ (4)
Here $|S_{fi}^{(\text{pw})}|^{2}$ defines the emission rate in momentum space where the phase [34, 35, 36] $\zeta_{fi}$ does not contribute to the probability, although it is non-vanishing even in the lowest order of the perturbation theory (Sec.5 in Supplementary). The result of the calculations is (Sec.3 in Supplementary)
@@
where $G(t^{\prime})>0$ is not important for now, the momentum conservation ${\bf p}={\bf p}^{\prime}+{\bf k}$ is implied, and
$\displaystyle{\bm{R}}={\bf r}-{\bm{u}}_{p}t+(\partial_{{\bf p}}+\partial_{{\bf k}})\zeta_{fi}({\bf p},\lambda_{e},{\bf k},\lambda_{\gamma}).$ (6)
Here ${\bm{u}}_{p}={\bf p}/\varepsilon({\bf p}),\,\varepsilon({\bf p})=\sqrt{{\bf p}^{2}+m^{2}},\,{\bm{u}}_{k}={\bf k}/(n|{\bf k}|),|{\bm{u}}_{k}|=1/n$, $F(t^{\prime})\propto\arctan t^{\prime}/t_{d}$ contains a Gouy phase of the photon connected to its spreading with time $t^{\prime}$, and $t_{d}$ is a diffraction time (see below).
The principal behavior of the Wigner function (5), which is not everywhere positive, is governed by the ratio
$\displaystyle\frac{{\displaystyle R^{2}}}{{\displaystyle R_{\text{eff}}^{2}(t^{\prime})}}=\frac{{\displaystyle 1}}{{\displaystyle\sigma_{x}^{2}(t^{\prime})}}\Bigg{(}\underbrace{\frac{{\displaystyle[{\bm{R}}\times({\bm{u}}_{p}-{\bm{u}}_{k})]^{2}}}{{\displaystyle({\bm{u}}_{p}-{\bm{u}}_{k})^{2}}}}_{\text{finite at $t^{\prime}=0$}}+\text{terms vanishing at}\,t^{\prime}=0\Bigg{)},$ (7)
where $\sigma_{x}^{2}(t^{\prime})=\sigma^{-2}\left(1+(t^{\prime}/t_{d})^{2}\right)$ is an rms size of the electron packet. When the condition of ChR is met, $u_{p}>u_{k}$, the vector ${\bm{u}}_{k}-{\bm{u}}_{p}$ is directed backwards with respect to the electron velocity ${\bm{u}}_{p}$, and dependence of the Wigner function on ${\bm{R}}$ at small $t^{\prime}$ vanishes along ${\bm{u}}_{k}-{\bm{u}}_{p}$, defining the Mach cone with an angle
@@
If the electron is detected in a plane-wave state, scattered at the angles $\theta^{\prime},\phi^{\prime}$, the radius $R_{\text{eff}}(t^{\prime})$ depends on the difference $\phi_{R}-\phi$ between the azimuthal angle of ${\bm{R}}$ and that of ${\bf k}$, so it is anisotropic. The azimuthal symmetry of the Mach cone is restored when the electron is not detected and we integrate Eq.(5) over ${\bf p}^{\prime}$.
@@
Figure 2: The effective correlation radius from Eq.(7) divided by the distance $u_{p}t^{\prime}$ traveled by the electron during the time $t^{\prime}$. Left: $\beta=0.7\,(\gamma=1.4),n=1.5,\,\theta_{\text{Ch.cl.}}=\arccos 1/u_{p}n\approx 17.8$ deg, $\theta_{\text{Mach}}\approx 107.8$ deg. Right: $\beta=0.9999\,(\gamma=70.7),n=1.33,\,\theta_{\text{Ch.cl.}}\approx 41.2$ deg, $\theta_{\text{Mach}}\approx 131.2$ deg. Nearby the Mach angle, space-time dependence of the Wigner function quickly vanishes within the correlation radius $R<R_{\text{eff}}(t^{\prime})$.
@@
Figure 3: The spreading time (9) of the photon field measured in picoseconds (a. and d.) and its inverse (b.). The sharp maxima are nearby the angle $\theta_{\text{Ch.cl.}}=\arccos 1/u_{p}n$. The Cherenkov condition is not met for the black line in a.), which is why the photon field quickly spreads during hundreds of attoseconds. The right upper panel shows two points (10), between which the spreading time turns negative due to the quantum recoil ($\omega/\varepsilon\neq 0$, see Eq.(11)).
@@
At large $t^{\prime}$, we find $R_{\text{eff}}^{2}(t^{\prime})\propto\sigma^{2}_{x}(0)\,t^{\prime 2}/t_{d}^{2}$ and when $R_{\text{eff}}(t^{\prime})\gg R$ dependence of the Wigner function on ${\bf r}$ and $t$ vanishes, whereas in the other limiting case, $R\gg R_{\text{eff}}(t^{\prime})$, it is exponentially suppressed. So, an effective region of space-time correlation is when $t^{\prime}\lesssim t_{d}$, and $R\sim R_{\text{eff}}(t^{\prime})$ is where it is most pronounced, which is why one can call $R_{\text{eff}}(t^{\prime})$ the correlation radius. At $t^{\prime}\gg t_{d}$, both the electron packet and the photon field spread and so there is no longer space-time correlation within the region $R<R_{\text{eff}}(t^{\prime})$, which is a hallmark of the wave zone. We show in Fig.2 that the correlation radius is orders of magnitude smaller than the distance $u_{p}t^{\prime}$ traveled by the electron for all the angles $\theta_{R}$ of ${\bm{R}}$, except for the Mach angle, $\theta_{R}\approx\theta_{\text{Mach}}$ (shown in Fig.3 c.)).
@@
$\displaystyle t_{d}=\frac{2}{\sigma^{2}}\frac{({\bm{u}}_{p}-{\bm{u}}_{k})^{2}}{\left(1/\omega n^{2}-\varepsilon^{-1}\right)({\bm{u}}_{p}-{\bm{u}}_{k})^{2}+\left(\varepsilon^{-1}-\omega^{-1}\right)\left[{\bm{u}}_{p}\times{\bm{u}}_{k}\right]^{2}}.$ (9)
where $({\bm{u}}_{p}-{\bm{u}}_{k})^{2}=n^{-2}+u_{p}^{2}-2u_{p}\cos\theta/n,\ \left[{\bm{u}}_{p}\times{\bm{u}}_{k}\right]^{2}=u_{p}^{2}\sin^{2}\theta/n^{2}$ with the $z$ axis directed along the electron partial momentum ${\bf p}$. One can also define the formation length of radiation as $L_{f}=u_{p}|t_{d}|$, which turns to infinity at the Cherenkov angle in the classical Tamm problem [37, 11, 12, 38].
The diffraction time and the formation length have an extremum either at the angle $\cos\theta_{\text{Ch.cl.}}=1/u_{p}n<1$, when the Cherenkov condition $u_{p}>1/n$ is met, or at $\cos\theta=u_{p}n<1$ otherwise. Importantly, the Wigner function and the energy density stay finite in the latter case, but the photon field rapidly spreads (see the black line in Fig.3 a.)). Along with the extremum, the time $t_{d}$ has two points in which its denominator vanishes (see Fig.3 b.)),

### v3

@@ Abstract
Cherenkov radiation of charged particles moving with superluminal velocities in transparent media is a well-studied phenomenon with a plethora of applications. Its microscopic origins can be traced to the polarization of atomic shells, characterized by time scales in the subfemtosecond range — dynamics that eludes conventional macroscopic treatment. Here we present a theoretical framework for probing the intrinsic dynamics of Cherenkov radiation, unveiling quantum features absent in classical realm and even in a fully quantum theory in momentum space. These features include a finite formation length and spreading time of the photon, the latter becoming negative nearby the Cherenkov angle, a finite flash duration tied to the size of the electron packet, along with a shift in the photon arrival time that can be either positive or negative and necessitates going beyond the far-field approximation. The calculated time scales lie in the attosecond range for the relevant parameters, thus linking this macroscopic phenomenon back to its atomic origins. Finally, we propose that by measuring the duration of the Cherenkov flash one can in principle retrieve the length of the emitting packet, deepening our understanding of quantum coherence effects in photon emission.
@@ Introduction
Cherenkov radiation (ChR) by charged particles in media [1, 2, 3, 4, 5, 6, 7, 8, 9, 10] is the simplest example of a wide range of phenomena embracing transition radiation, diffraction, Smith-Purcell radiation, and other mechanisms of photon emission [11, 12, 13, 14]. Their common microscopic origin is atomic bremsstrahlung [15] due to dynamic polarization of atomic shells by the field of the charge, and the characteristic time scales are femto- and attoseconds, typical for the AC Stark effect [16]. Along with numerous applications of classical ChR for particle detection in acceleration experiments and cosmic rays studies [17], neutrino telescopes [18, 19], gamma-ray astronomy [20, 21], and other fields [10], it has also recently attracted attention of the biomedical community as a new tool for molecular imaging and therapy in cancer treatment [22, 23, 24].
A quantum theory of ChR in momentum space was presented in 1940 by Ginzburg [4] and Sokolov [5] who predicted a correction to the classical Tamm-Frank result due to quantum recoil, which is negligibly small for the overwhelming majority of applications of ChR. Interest in inherently quantum features of ChR was revived in 2016 [25, 26, 27] after vortex electrons with quantized orbital angular momentum projection were generated at electron microscopes [28, 29]. Some flaws in the initial analysis of Ref.[25] were corrected in Ref.[26]. More recently, it has been argued that the spatiotemporal features of ChR can be connected to the spatial coherence of the emitting charged particle [30], whereas generalized measurements of either the final electron or the photon can lead to the generation of a wave packet of the other particle with the needed properties [31].
The spatiotemporal characteristics of radiation in the pre-wave zone are neccesary for applications of such close relatives of ChR as transition, diffraction, and Smith-Purcell radiation, for instance, in bunch-length measurements at accelerators [32, 12]. The classical theory of the Cherenkov wake fields in the near-field zone has also been developed for beam position monitors and radiation sources at accelerators and free-electron lasers [33, 34, 35] and even alternative concepts of the Cherenkov wakefield acceleration have been proposed [36]. However, no quantum theory of ChR and its generalizations exists so far beyond the far-field approximation. The latter is tightly linked to the conventional momentum-space approach, whereas a spatiotemporal analysis of the photon field at a finite distance from the emitting particle necessitates working in real space and time, and that is always tricky in relativistic quantum theory. Although the flash duration of ChR was first estimated by Frank as early as in 1956 from classical considerations [8, 9], the microscopic atomic dynamics in the formation of the ChR field remains hidden even in the fully quantum treatment of ChR in momentum space.
@@
Here we point out that one can access atomic time scales in ChR and in its generalizations by using a quantum theory in phase space where – similar to quantum optics [37] – we employ a Wigner function to characterize the emitted photon field. We demonstrate how to probe the field in the pre-wave – or formation – zone [32, 12, 38] in which the partial waves interfere, the Cherenkov cone is not formed yet, and the emitted energy propagates in real space and time as a spreading wave packet. We find that our theory establishes a link between the evolving coherence length of the electron packet, the Cherenkov flash duration, and a quantum temporal delay that the photon experiences in medium and that can only be quantitatively studied beyond the far-field approximation. This delay falls within the attosecond range, a time scale characteristic of atomic excitation processes explored in attosecond spectroscopy and metrology [39], and it can be either positive or negative, that is, the photon wave packet can reach the detector on average later or sooner than one can expect from the classical far-field considerations. This temporal detuning also shares some similarities with tunneling-related time effects, such as those observed in electron ionization [40] and emission [41]. The quantum shift in the photon arrival time, coupled with the finite flash duration and the spreading dynamics, unveils intricate coherence properties that enable refined temporal control in quantum emission processes, enriching the landscape of quantum optics, ultrafast physics, and of applications of ChR and related phenomena.
Figure 1: Spatiotemporal dynamics of the Cherenkov radiation generated by a spreading electron wavepacket. $\sigma_{x}$ stands for the initial size of an electron packet (blue), $e$, and $n$ is the refractive index of the medium. The radiation process begins with the fundamental quantum electrodynamics interaction $e\rightarrow e^{\prime}+\gamma$, during which the photon field (pink), $\gamma$, gradually separates from the electron packet within the formation zone and spreads. The electron is detected with a momentum ${\bf p}^{\prime}=|{\bf p}^{\prime}|\{\sin\theta^{\prime}\cos\phi^{\prime},\sin\theta^{\prime}\sin\phi^{\prime},\cos\theta^{\prime}\}$, and the Cherenkov cone is formed in the far-field only if the electron remains undetected. In phase-space picture, the photon with the momentum ${\bf k}=n\omega\{\sin\theta\cos\phi,\sin\theta\sin\phi,\cos\theta\}$ can be detected at a finite distance, meaning it does not necessarily reach the far field.
Results
@@
The system of units $\hbar=c=1$ is used, the electron velocity is $\beta=u_{p}/c\equiv u_{p}$, $m=0.511$ MeV is the electron mass, $1/m\equiv\hbar/mc=3.86\times 10^{-11}$ cm is the reduced Compton wavelength, and the corresponding timescale is $t_{\text{C}}=1/m\equiv\hbar/mc^{2}\approx 1.3\times 10^{-21}\,\text{s}$. When an electron emits a photon, the two-particle state within the first order of the perturbation theory in quantum electrodynamics is $|e^{\prime},\gamma\rangle=(\hat{1}+\hat{S}^{(1)})|\text{in}\rangle$, where $|\text{in}\rangle=|e_{\text{in}}\rangle\otimes|0_{\gamma}\rangle$ and $\hat{S}^{(1)}=-ie\int d^{4}x\,\hat{j}^{\mu}(x)\hat{A}_{\mu}(x)$ [42]. Fig.1 illustrates how the photon field gradually becomes spatially separated from the electron packet, spreads, and eventually turns into a plane wave propagating at the angles $\theta,\phi$. If we detect the electron in a plane-wave state $\langle{\bf p}^{\prime},\lambda^{\prime}|$, $\lambda^{\prime}=\pm 1/2$, the state of the photon becomes
$\displaystyle|\gamma\rangle=\langle{\bf p}^{\prime},\lambda^{\prime}|e_{\text{in}}\rangle|0_{\gamma}\rangle+\sum\limits_{\lambda_{\gamma}}\int\frac{d^{3}k}{(2\pi)^{3}}\,|{\bf k},\lambda_{\gamma}\rangle\,S_{\text{fi}}.$ (1)
The second term here is a coherent superposition of plane waves with the momenta ${\bf k}$ and the helicity $\lambda_{\gamma}=\pm 1$ and a matrix element is $S_{\text{fi}}=\langle{\bf k},\lambda_{\gamma};{\bf p}^{\prime},\lambda^{\prime}|\hat{S}^{(1)}|\text{in}\rangle$. We treat the incoming electron as a Gaussian packet with a wave function $f_{\text{e}}^{(\text{in})}({\bf p},\lambda)=\langle{\bf p},\lambda|e_{\text{in}}\rangle$, a mean momentum $\langle{\bf p}\rangle$, an uncertainty $\sigma\ll m$, which means that the rms size – also called coherence length – of the packet is much larger than the Compton wavelength, $\sigma_{x}=1/\sigma\gg 1/m$. For the moment, we take a simplified model with the packet, spherically symmetric in the laboratory frame, $\sigma_{\perp}=\sigma_{||}\equiv\sigma_{x}=1/\sigma$.
Let us define the energy density of the photon field in real space and time, an electric part of which is (Supplementary Notes 1 and 2)
$\displaystyle\frac{{\displaystyle 1}}{{\displaystyle 8\pi}}\langle\gamma|\hat{\bm{E}}^{2}({\bf r},t)|\gamma\rangle\to\frac{{\displaystyle 1}}{{\displaystyle 4\pi}}\left|\langle 0|\hat{\bm{E}}({\bf r},t)|\gamma\rangle\right|^{2}=\int\frac{d^{3}k}{(2\pi)^{3}}\,\mathcal{W}({\bf r},{\bf k},t),$ (2)
where we have subtracted a contribution of the vacuum energy and $\hat{\bm{E}}({\bf r},t)$ is a secondary-quantized electric field operator [43]. A contribution of the magnetic field can be written in a similar way. Here
@@
is a Wigner function of the photon, and $\bm{E}_{\lambda_{\gamma}}({\bf k})=\frac{i\omega\sqrt{4\pi}}{\sqrt{2\omega n^{2}}}\,{\bm{e}}_{{\bf k}\lambda_{\gamma}}\sum\limits_{\lambda}\int\frac{d^{3}p}{(2\pi)^{3}}\,f_{\text{e}}^{(\text{in})}({\bf p},\lambda)S_{\text{fi}}^{(\text{pw})},\,{\bm{e}}_{{\bf k}\lambda_{\gamma}}\cdot{\bf k}=0,$ is a positive-frequency component of its electric field.
According to the standard interpretation [43], the energy density Eq. (2) defines probability of detecting a photon in a region of space centered at the point ${\bf r}$ at the moment of time $t$. Clearly, the second marginal distribution $\int d^{3}x\,\mathcal{W}({\bf r},{\bf k},t)\propto|S_{\text{fi}}^{(\text{pw})}|^{2}$ yields probability to detect a photon with the frequency $\omega$ and a wave vector ${\bf k},\,|{\bf k}|=n(\omega)\,\omega$, the standard result of the quantum theory in momentum space (see [26]). Therefore, it is this Wigner function Eq. (3) in phase space that contains all the information on spatial distribution of the photon energy density at a given distance ${\bf r}$ – also in the near-field zone – and on its dynamics. Consequently, emission takes place in the pre-wave zone even if the condition of ChR is not met, but the waves do not constructively interfere to form a cone in the far field.
We calculate the Wigner function in the paraxial approximation, $\sigma\ll m$, in a medium with weak dispersion, $\frac{\omega}{n(\omega)}\frac{dn(\omega)}{d\omega}\ll 1$, and represent the tree-level amplitude [42, 26] as follows:
$S_{\text{fi}}^{(\text{pw})}=|S_{\text{fi}}^{(\text{pw})}|\,e^{i\zeta_{\text{fi}}}.$ (4)
Here $|S_{\text{fi}}^{(\text{pw})}|^{2}$ defines the emission rate in momentum space where the phase [28, 44, 29] $\zeta_{\text{fi}}$ does not contribute to the probability, although it is non-vanishing even in the lowest order of the perturbation theory (see Supplementary Note 5). The result of the calculations is (see Supplementary Note 3)
@@
where $G(t^{\prime})>0$ is the function particular form of which is of no concern for the following discussion, the momentum conservation ${\bf p}={\bf p}^{\prime}+{\bf k}$ is implied, and
$\displaystyle{\bf R}={\bf r}-{\bm{u}}_{p}t+(\partial_{{\bf p}}+\partial_{{\bf k}})\zeta_{\text{fi}}({\bf p},\lambda_{e},{\bf k},\lambda_{\gamma}).$ (6)
Here ${\bm{u}}_{p}={\bf p}/\varepsilon({\bf p}),\,\varepsilon({\bf p})=\sqrt{{\bf p}^{2}+m^{2}},\,{\bm{u}}_{k}={\bf k}/(n|{\bf k}|),|{\bm{u}}_{k}|=1/n$, $\partial_{{\bf p}}=\partial/\partial{\bf p}$, $F(t^{\prime})\propto\arctan t^{\prime}/t_{\text{d}}$ contains a Gouy phase of the photon connected to its spreading with time $t^{\prime}$, and $t_{\text{d}}$ is a diffraction time (see below).
The spatio-temporal dependence of the Wigner function Eq. (5), which is not everywhere positive even in the paraxial approximation, is governed by the ratio
$\displaystyle\frac{{\displaystyle R^{2}}}{{\displaystyle R_{\text{eff}}^{2}(t^{\prime})}}=\frac{{\displaystyle 1}}{{\displaystyle\sigma_{x}^{2}(t^{\prime})}}\Bigg{(}\underbrace{\frac{{\displaystyle[{\bf R}\times({\bm{u}}_{p}-{\bm{u}}_{k})]^{2}}}{{\displaystyle({\bm{u}}_{p}-{\bm{u}}_{k})^{2}}}}_{\text{finite at $t^{\prime}=0$}}+\text{terms vanishing at}\,t^{\prime}=0\Bigg{)},$ (7)
where $\sigma_{x}^{2}(t^{\prime})=\sigma^{-2}\left(1+(t^{\prime}/t_{\text{d}})^{2}\right)$ is an rms size of the electron packet. When the condition of ChR is met, $u_{p}>u_{k}$, the vector ${\bm{u}}_{k}-{\bm{u}}_{p}$ is directed backwards with respect to the electron velocity ${\bm{u}}_{p}$, and dependence of the Wigner function on ${\bf R}$ at small $t^{\prime}$ vanishes along ${\bm{u}}_{k}-{\bm{u}}_{p}$, defining the Mach cone with an angle
@@
If the electron is detected in a plane-wave state, scattered at the angles $\theta^{\prime},\phi^{\prime}$, the radius $R_{\text{eff}}(t^{\prime})$ depends on the difference $\phi_{R}-\phi$ between the azimuthal angle of ${\bf R}$ and that of ${\bf k}$, so it is anisotropic. The azimuthal symmetry of the Mach cone is restored when the electron is not detected and we integrate Eq.(5) over ${\bf p}^{\prime}$.
@@
Figure 2: The effective correlation radius of the Cherenkov radiation. The ratio $R_{\rm{eff}}$ to the distance $u_{p}t^{\prime}$ is given for the electron traveling times $t^{\prime}=10^{7}\,t_{\text{C}}$ (solid black line), $t^{\prime}=10^{10}\,t_{\text{C}}$ (dotted red line) and $t^{\prime}=10^{14}\,t_{\text{C}}$ (dot-dashed green line). At panel (a): $\beta=0.7\,(\gamma=1.4),n=1.5,\,\theta=\theta_{\text{Ch.cl.}}=\arccos(1/u_{p}n)\approx 17.8^{\circ},\,\omega=10^{-6}m,\,\sigma=10^{-5}m,\,\phi_{R}-\phi=0$ deg, $\theta_{\text{Mach}}\approx 107.8^{\circ}$. At panel (b): $\beta=0.9999\,(\gamma=70.7),n=1.33,\,\theta=\theta_{\text{Ch.cl.}}\approx 41.2^{\circ},\omega=10^{-5}m,\,\sigma=10^{-4}m,\,\phi_{R}-\phi=0$ deg, $\theta_{\text{Mach}}\approx 131.2^{\circ}$. Nearby the Mach angle $\theta_{\text{Mach}}$, space-time dependence of the Wigner function quickly vanishes within the correlation radius $R<R_{\text{eff}}(t^{\prime})$, which is a hallmark of the wave zone.
@@
Figure 3: The spreading time $t_{\text{d}}$ of the photon field, angular spans $\Delta\theta_{\infty}$ defining negative areas for $t_{\text{d}}$ and the Mach cone of the emission. Panel (a): We take $\omega=0.51$ eV, $\sigma=10^{-5}$ and the packet width $\sigma_{x}=1/\sigma\approx 38$ nm. The spreading time $t_{\text{d}}$ (in picoseconds) from Eq. (9) is displayed for the electron velocity $\beta=0.65$ and refractive index $n=1.46$ (solid black), $\beta=0.7$ and $n=1.46$ (dashed blue), $\beta=0.9999$ and $n=1.05$ (dotted green) and $\beta=0.9999$ and $n=1.3$ (dot-dashed red line). The sharp maxima are nearby the angle $\theta_{\text{Ch.cl.}}=\arccos 1/u_{p}n$ as illustrated by the arrows. The Cherenkov condition is not met for the black line, which is why the photon field quickly spreads during hundreds of attoseconds. Panel (b): The inverse spreading time is plotted for parameters $n=1.5,\,\beta=0.9,\,\sigma=10^{-4}$m and $\omega=51.1$ eV (solid black), $\omega=5.1$ eV (dashed blue) and $\omega=0.51$ eV (dotted green line). In the angular span of a size $\Delta\theta_{\infty}$ between two points given by Eq. (10) the spreading time turns negative due to the quantum recoil ($\omega/\varepsilon\neq 0$, see Eq. (11)). Panel (c): A close-up picture of the spreading time $t_{\text{d}}$ for parameters $n=1.5,\,\beta=0.999,\,\omega=5.1$ eV, $\sigma=10^{-4}$m, $\theta_{\text{Ch.cl}}\approx 48.1$ deg. Panel (d): Formation of the Mach cone of the electron emission. The electron moves with velocity ${\bm{u}}_{p}$ exceeding the photon velocity ${\bm{u}}_{k}$. The photon field produces the Mach cone with an angle $\theta_{\text{Mach}}$ according to Eq.(8).
@@
At large $t^{\prime}$, we find $R_{\text{eff}}^{2}(t^{\prime})\propto\sigma^{2}_{x}(0)\,t^{\prime 2}/t_{\text{d}}^{2}$ and when $R_{\text{eff}}(t^{\prime})\gg R$ dependence of the Wigner function on ${\bf r}$ and $t$ vanishes, similar to the classical wave zone. In the other limiting case, $R\gg R_{\text{eff}}(t^{\prime})$, the integrand in Eq.(5) is exponentially suppressed. So, an effective region where space-time correlation exists is when $t^{\prime}\lesssim t_{\text{d}}$, and $R\sim R_{\text{eff}}(t^{\prime})$ is where it is most pronounced, which is why one can call $R_{\text{eff}}(t^{\prime})$ the correlation radius. At $t^{\prime}\gg t_{\text{d}}$, both the electron packet and the photon field spread and so there is no longer space-time correlation within the region $R<R_{\text{eff}}(t^{\prime})$, which is a hallmark of the wave zone. We show in Fig.2 that the correlation radius is orders of magnitude smaller than the distance $u_{p}t^{\prime}$ traveled by the electron during the time interval $t^{\prime}$ for all the angles $\theta_{R}$ of ${\bf R}$, except for the Mach angle, $\theta_{R}\approx\theta_{\text{Mach}}$ (shown in Fig.3 d).
@@
$\displaystyle t_{\text{d}}=\frac{2}{\sigma^{2}}\frac{({\bm{u}}_{p}-{\bm{u}}_{k})^{2}}{\left(1/\omega n^{2}-\varepsilon^{-1}\right)({\bm{u}}_{p}-{\bm{u}}_{k})^{2}+\left(\varepsilon^{-1}-\omega^{-1}\right)\left[{\bm{u}}_{p}\times{\bm{u}}_{k}\right]^{2}}.$ (9)
where $({\bm{u}}_{p}-{\bm{u}}_{k})^{2}=n^{-2}+u_{p}^{2}-2u_{p}\cos\theta/n,\ \left[{\bm{u}}_{p}\times{\bm{u}}_{k}\right]^{2}=u_{p}^{2}\sin^{2}\theta/n^{2}$ with the $z$ axis directed along the electron partial momentum ${\bf p}$. One can also define the formation length of radiation as $L_{\text{f}}=u_{p}|t_{\text{d}}|$, which turns to infinity at the Cherenkov angle in the classical Tamm problem [45, 11, 12, 46].
The diffraction time and the formation length have an extremum either at the angle $\cos\theta_{\text{Ch.cl.}}=1/u_{p}n<1$, when the Cherenkov condition $u_{p}>1/n$ is met, or at $\cos\theta=u_{p}n<1$ otherwise. The Wigner function and the energy density stay finite in the latter case, but the photon field rapidly spreads (see the black line in Fig.3 a). Along with the extremum, the time $t_{\text{d}}$ has two points in which its denominator vanishes (see Fig.3 b),